OpenStudy (anonymous):

Anusha brings a $20 bill to the store. She purchases 4 pairs of socks for $2.25 each and a new belt for $7.50. How much change will Anusha receive when she pays for her clothes? $3.50 $7.50 $9.00 $16.50

OpenStudy (lieutenantgeneral):

2.25 * 4 = 9 20 - 9 = 11 11 - 7.50 = 3.50 Answer: $3.50